( 3xy + 3y - 4)dx + (x + 1)2dy = 0
1
Expert's answer
2020-11-18T19:28:03-0500

Let us solve the equation "( 3xy + 3y - 4)dx + (x + 1)^2dy = 0":


"3xy + 3y - 4 + (x + 1)^2y' = 0"


"3(x + 1)y + (x + 1)^2y' = 4"


Multiply both parts of the differential equation by "x+1:"


"3(x + 1)^2y + (x + 1)^3y' = 4(x+1)"


"((x + 1)^3y)' = 4x+4"


Therefore, we have the following solution:


"(x + 1)^3y = 2x^2+4x+C"